An object has a launch velocity of 30.0 m/s at launch angle of 30.00 relative to horizontal. Find the a) x and y component of initial velocity, b) time to reach the peak c) total time of flight, d) vertical height to reach the peak, and e) horizontal distance. 


1
Expert's answer
2022-03-14T09:45:42-0400

a)

"v_{0x}=v_0\cos\theta\\
=30.0\cos30.0^\circ=26.0\:\rm m/s"

"v_{0y}=v_0\sin\theta\\
=30.0\sin30.0^\circ=15.0\:\rm m/s"

b)

"t_1=\frac{v_{0y}}{g}=\frac{26.0}{9.81}=2.65\:\rm s"

c)

"t_2=\frac{2v_{0y}}{g}=\frac{2*26.0}{9.81}=5.30\:\rm s"

d)

"h_{\max}=\frac{v_{0y}^2}{2g}=\frac{26.0^2}{2*9.81}=34.5\:\rm m"

e)

"R=v_{0x}t_2=15.0*5.30=79.5\:\rm m"